Hi. The instructions say "Use a table of values to graph the equation." Question is 4x + 4y = 2 Thanks!
Found 2 solutions by rfer, dpaulin:
Answer by rfer(16322)

create an input/output table the graph the results.
4x+4y=2
4(0)+4y=2
y=1/2
(0,1/2)
------------
4x+4(0)=2
4x=2
x=1/2
(1/2,0)
----------------
solve at least one more then graph

Answer by dpaulin(12)